OpenStudy (anonymous):

1.13 area of a Square.. 1.Find the area of a square if a side is a.45 cm b.9.25 m c. 6 1/2 dm d. 35 2/3 cm ______ ________ _________ _________ 2. Find the side of a square whose area is 900 sq.cm. _____________________ 3. If one side of a square tablecloth is 120 cm, what is the area of the table if the cloth hangs 20 cm from the table???? ___________ 4. a side of a square rice field is 350 m. How many hectares is the rice field? if one hectare costs 500,000, How much does the field cost?? PLSS HELP ME..

OpenStudy (anonymous):

1.13 area of a Square.. 1.Find the area of a square if a side is a.45 cm b.9.25 m c. 6 1/2 dm d. 35 2/3 cm ______ ________ _________ _________ 2. Find the side of a square whose area is 900 sq.cm. _____________________ 3. If one side of a square tablecloth is 120 cm, what is the area of the table if the cloth hangs 20 cm from the table???? ___________ 4. a side of a square rice field is 350 m. How many hectares is the rice field? if one hectare costs 500,000, How much does the field cost?? PLSS HELP ME..

OpenStudy (anonymous):

why are you posting so many questions?

OpenStudy (anonymous):

Nothing...

OpenStudy (anonymous):

The area of a square is one of its sides squared....for example A square that has a side that is 45 cm long will have an area of \[45^2\]

OpenStudy (anonymous):

Can you do the rest of question 1 now?

OpenStudy (anonymous):

Hmmm.

OpenStudy (anonymous):

What i answer to number 2?

OpenStudy (anonymous):

square root th egiven value

OpenStudy (anonymous):

So what answer in number 2??

OpenStudy (anonymous):

squareroot(900)

OpenStudy (anonymous):

\[\huge \color{DarkGreen} {Area = (Side)^2}\]

OpenStudy (anonymous):

\[\huge \color{red}{Side = \sqrt{Area}}\]

OpenStudy (anonymous):

In number 3 what is the answer and number 4/>>.????

OpenStudy (anonymous):

In 3rd, first you have to decrease 20 from 120 120 - 20 = 100 Now this is the side.. now find area.. formula I have written above... Use and tell me what you will get..

OpenStudy (anonymous):

Doing or not @D.Wade ??

OpenStudy (anonymous):

I do not really know. please I'm weak in math

OpenStudy (anonymous):

What is the square of 100??

OpenStudy (anonymous):

Okay.. tell me what you will get when you multiply 100 with 100??

OpenStudy (anonymous):

\[\huge \color{green}{100 \times 100} = ?\;?\;?\]

OpenStudy (anonymous):

100 multiply 100 = 10,000

OpenStudy (anonymous):

I think we are going wrong.. Tell me what is: \[\huge \color{red}{120 \times 120 = ?\;?}\]

OpenStudy (anonymous):

The square root of 100 is : ± 10. This is because 10 x 10 = 100 and -10 x -10 = 100.

OpenStudy (anonymous):

I am asking you different... 120 X 120 = ???

OpenStudy (anonymous):

14,400

OpenStudy (anonymous):

And what is : \[\huge 20 \times 20 = ??\]

OpenStudy (anonymous):

400

OpenStudy (anonymous):

400

OpenStudy (anonymous):

Now subtract them: \[\huge 14400 - 400 = ??\]

OpenStudy (anonymous):

14,000

OpenStudy (anonymous):

Yes dear this is your answer for number 3..

OpenStudy (anonymous):

In number 2 what is answer ??

OpenStudy (anonymous):

What is the square root of 900??

OpenStudy (anonymous):

question number 2... Find the side of square whose area is 900 sq.cm.

OpenStudy (anonymous):

I am asking you: \[\huge \sqrt{900} = ??\]

OpenStudy (anonymous):

w8...

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

30.?>?

OpenStudy (anonymous):

yes this will be the side..

OpenStudy (anonymous):

so 30 is the answer in number 2.

OpenStudy (anonymous):

In number 1. letter b and c and d what is the answer??

OpenStudy (anonymous):

Take the square of each and every term you will have the answer..

OpenStudy (anonymous):

In number 4 what is the answer??

OpenStudy (ganpat):

For the 4th one.. 1 m² = 0.0001 hec so we have Area of = 350 * 350 m square... = (350 * 350 ) * 0.0001 = 12.25 Hec .. Now, 1 hector cost 5,00,000.. So 12.25 will cost 5,00,000 * 12.25 = 6125000..

OpenStudy (anonymous):

What is the answer in number 4? 6125000??